# OR-LLM-Agent: Automating Modeling and Solving of Operations Research Optimization Problems with Reasoning LLM

### v2

@@ Introduction
Due to known inaccuracies in reference answers in NL4OPT, MAMO-Easy, MAMO-Complex, and IndustryOR, we report only code error rates for these datasets. On BWOR, both code and mathematical model accuracy are evaluated. The results are shown in Table 2 and Figure 12.
Model (%) IndustryOR ComplexLP EasyLP NL4OPT BWOR
OR-LLM-Agent-GPT-o3 1.00 0.00 0.00 0.00 1.22
GPT-o3 4.00 4.74 1.07 0.41 3.66
OR-LLM-Agent-Gemini 2.5 Pro 0.00 0.00 0.00 0.00 0.00
Gemini 2.5 Pro 6.00 3.32 1.07 0.41 2.44
OR-LLM-Agent-DeepSeek-R1 3.00 0.82 0.00
DeepSeek-R1 13.00 3.32 1.38 1.22 4.88
OR-LLM-Agent-GPT-4o 1.00 0.00 0.15 2.45 2.44
GPT-4o 17.00 1.90 2.30 6.53 9.76
OR-LLM-Agent-Gemini 2.0 Flash 2.00 0.00 0.00 0.00 0.00
Gemini 2.0 Flash 8.00 3.79 0.15 0.00 4.88
OR-LLM-Agent-DeepSeek-V3 0.00 0.47 0.15 0.00 0.00
DeepSeek-V3 12.00 2.37 1.07 2.86 13.41
Table 2: Code error rates of LLMs across OR datasets.
Figure 12: Effect of OR-LLM-Agent on mathematical model accuracy within runnable code.
OR-LLM-Agent reduces the average code error rate by 3.74% across all datasets. On BWOR, it increases the mathematical model accuracy by 5.77% on average. For DeepSeek-V3, although the code error rate decreases, the mathematical model accuracy within the runnable code drops by 2.86%. This may be attributed to OR-LLM-Agent’s stronger ability to repair code than to improve the underlying mathematical modeling when used with DeepSeek-V3.
In summary, these results suggest that OR-LLM-Agent enhances code executability and, in most cases, improves mathematical modeling accuracy in runnable code.
@@ Conclusion
This paper proposes OR-LLM-Agent, an AI agent framework for solving OR problems, built upon reasoning LLMs. The proposed approach enables a fully automated problem-solving pipeline without requiring any additional fine-tuning or complex prompt engineering. OR-LLM-Agent emulates the human problem-solving process by decomposing OR tasks into three sequential subtasks: mathematical modeling, coding, and debugging. Each subtask is handled independently by a dedicated agent, enabling targeted reasoning and improving overall performance. To evaluate LLM performance on OR tasks, we construct BWOR, a high-quality OR dataset. Experimental analysis suggests that NL4OPT, MAMO, and IndustryOR datasets may contain issues that limit their reliability for evaluating LLM performance on OR tasks. Compared to existing datasets, BWOR more effectively differentiates models of varying capabilities. Experimental results show OR-LLM-Agent significantly enhances the performance of LLMs on OR tasks while also reducing error rates during problem solving. These results demonstrate the effectiveness of combining reasoning LLMs with task decomposition for solving OR problems.

### v3

@@ Introduction
On NL4OPT, MAMO-Easy, MAMO-Complex, and IndustryOR datasets, we report code error rates. On BWOR, both code and mathematical model accuracy are evaluated. The results are shown in Table 2 and Table 3. We evaluate two groups: (a) OR-LLM-Agent with different LLMs as sub-agents, and (b) reasoning and non-reasoning LLMs.
Group Model (%) ACC
OR- OR-LLM-Agent(GPT-o3) 80.25
LLM- OR-LLM-Agent(Gemini 2.5 Pro) 80.49
Agent OR-LLM-Agent(DeepSeek-R1) 82.93
OR-LLM-Agent(GPT-4o) 53.75
OR-LLM-Agent(Gemini 2.0 Flash) 65.85
OR-LLM-Agent(DeepSeek-V3) 69.51
Mean 72.13
Reason- GPT-o3 78.48
ing and Gemini 2.5 Pro 73.75
non- DeepSeek-R1 76.92
reason- GPT-4o 44.59
ing Gemini 2.0 Flash 52.56
LLMs DeepSeek-V3 71.83
Mean 66.35
Gap OR-LLM-Agent -- Reasoning and 5.78
non-reasoning LLMs (72.13-66.25)
Table 3: Mathematical model accuracy of LLMs.
Compared to reasoning and non-reasoning LLMs, OR-LLM-Agent reduces the mean code error rate by 4.04%, dropping from 4.56% to 0.52%. On BWOR, it improves mean mathematical model accuracy by 5.78%, from 66.35% to 72.13%. These results suggest that using the OR-LLM-Agent framework enhances code executability and, in most cases, improves mathematical modeling accuracy.
@@ Conclusion
This paper proposes OR-LLM-Agent, an AI agent framework for solving OR problems, built upon reasoning LLMs. The proposed approach enables a fully automated problem-solving pipeline without requiring any additional fine-tuning or complex prompt engineering. OR-LLM-Agent emulates the human problem-solving process by decomposing OR tasks, enabling targeted reasoning and improving overall performance. To evaluate LLM performance on OR tasks, we construct BWOR, an OR benchmark dataset. Compared to existing datasets, BWOR more effectively differentiates models of varying capabilities. Experimental results show the OR-LLM-Agent framework significantly enhances the performance of LLMs on OR tasks while also reducing error rates during problem solving. These results demonstrate the effectiveness of combining reasoning LLMs with task decomposition for solving OR problems.